# Sequence Modeling for N-Agent Ad Hoc Teamwork

### v2

@@ 2 Preliminaries
The N-agent ad hoc teamwork (NAHT) problem is formalized in the framework of Decentralized Partially Observable Markov Decision Processes (Dec-POMDPs). A Dec-POMDP is described by the tuple $\langle M,S,\{O^{i}\}_{i=1}^{M},A,T,r,\gamma,T\rangle$. Here, $M$ denotes the number of agents, $S$ the global state space, $A=\prod_{i=1}^{M}A^{i}$ the joint action space, and $O^{i}$ the local observation space for agent $i$. Let the notation $\Delta(\bullet)$ denote a probability distribution over the space $\bullet$. The transition function $T:S\times A\rightarrow\Delta(S)$ governs the dynamics, and the reward function $r:S\times A\rightarrow\mathbb{R}$ provides a shared reward signal. In this work, we examine a centralized control scheme where each agent selects actions according to the stochastic policy $\pi^{i}:H^{i}\times A^{1}\times\cdots A^{i-1}\rightarrow\Delta(A^{i})$, where $H^{i}$ is the history of observations and actions available to agent $i$, and the policy conditions on the actions selected by agents $1$ to $i-1$ in some fixed, arbitrary ordering. The decision-making problem occurs under the horizon $T$.
The NAHT problem considers a controlled set of agents $C$, and an uncontrolled set $U$. The learning algorithm generates policies for agents in $C$, while uncontrolled agents in $U$ follow fixed, potentially unknown policies. Each episode begins with a team sampling process $X(U,C)$, consisting of two steps: first, the number of controlled agents $N<M$ is sampled, and second, a set of $N$ agents is drawn from the controlled set $C$, while $M-N$ agents are drawn from the uncontrolled set $U$. The NAHT learning problem considers a parameterized set of controlled agents, $C(\theta)$, and aims to maximize the expected return under varying teammates and team compositions: $\max_{\theta}\mathbb{E}_{\{\pi^{i}\}_{i=1}^{M}\sim X(U,C(\theta))}\left[\sum_{t=0}^{T}\gamma^{t}r_{t}\right].$
@@ 4.1 Experimental Setting
Following the procedure outlined in Wang et al. [12], we train five types of uncontrolled teams using five MARL algorithms: VDN, IPPO, MAPPO, IQL, and QMIX [9, 14, 10, 5]. For a particular episode with $M$ total agents, an uncontrolled team is first sampled uniformly at random from the set of five team types. Next, the number of controlled agents is sampled uniformly, $N\in[1,M-1]$. In the episode, $N$ controlled agents must coordinate with $M-N$ uncontrolled teammates of the sampled type.

### v3

@@ 2 Preliminaries
The N-agent ad hoc teamwork (NAHT) problem is formalized in the framework of Decentralized Partially Observable Markov Decision Processes (Dec-POMDPs). A Dec-POMDP is described by the tuple $\langle M,S,\{O^{i}\}_{i=1}^{M},A,T,r,\gamma,T\rangle$. Here, $M$ denotes the number of agents, $S$ the global state space, $A=\prod_{i=1}^{M}A^{i}$ the joint action space, and $O^{i}$ the local observation space for agent $i$. Let the notation $\Delta(\bullet)$ denote a probability distribution over the space $\bullet$. The transition function $T:S\times A\rightarrow\Delta(S)$ governs the dynamics, and the reward function $r:S\times A\rightarrow\mathbb{R}$ provides a shared reward signal. In this work, we examine a centralized control scheme where each agent selects actions according to the stochastic policy $\pi^{i}:H^{i}\times A^{1}\times\cdots A^{i-1}\rightarrow\Delta(A^{i})$, where $H^{i}$ is the history of observations and actions available to agent $i$, and the policy conditions on the actions selected by agents $1$ to i1i-1 in some fixed, arbitrary ordering. The decision-making problem occurs under the horizon $T$.
The NAHT problem considers a controlled set of agents $C$, and an uncontrolled set $U$. The learning algorithm generates policies for agents in $C$, while uncontrolled agents in $U$ follow fixed, potentially unknown policies. Each episode begins with a team sampling process $X(U,C)$, consisting of two steps: first, the number of controlled agents $N<M$ is sampled, and second, a set of $N$ agents is drawn from the controlled set $C$, while MNM-N agents are drawn from the uncontrolled set $U$. The NAHT learning problem considers a parameterized set of controlled agents, $C(\theta)$, and aims to maximize the expected return under varying teammates and team compositions: $\max_{\theta}\mathbb{E}_{\{\pi^{i}\}_{i=1}^{M}\sim X(U,C(\theta))}\left[\sum_{t=0}^{T}\gamma^{t}r_{t}\right].$
@@ 4.1 Experimental Setting
Following the procedure outlined in Wang et al. [12], we train five types of uncontrolled teams using five MARL algorithms: VDN, IPPO, MAPPO, IQL, and QMIX [9, 14, 10, 5]. For a particular episode with $M$ total agents, an uncontrolled team is first sampled uniformly at random from the set of five team types. Next, the number of controlled agents is sampled uniformly, $N\in[1,M-1]$. In the episode, $N$ controlled agents must coordinate with MNM-N uncontrolled teammates of the sampled type.